# Towards Generalizable AI-Generated Image Detection via Image-Adaptive Prompt Learning

### v2

@@ Abstract
In AI-generated image detection, current cutting-edge methods typically adapt pre-trained foundation models through partial-parameter fine-tuning. However, these approaches often struggle to generalize to forgeries from unseen generators, as the fine-tuned models capture only limited patterns from training data and fail to reflect the evolving traits of new ones. To overcome this limitation, we propose Image-Adaptive Prompt Learning (IAPL), a novel paradigm that dynamically adjusts the prompts fed into the encoder according to each input image, rather than fixing them after training. This design significantly enhances robustness and adaptability to diverse forged images. The dynamic prompts integrate conditional information with test-time adaptive tokens through a lightweight gated mechanism. The conditional information is produced by a Conditional Information Learner, which leverages CNN-based feature extractors to model both forgery-specific and image-specific conditions. The test-time adaptive tokens are optimized during inference on a single sample by enforcing prediction consistency across multiple views, ensuring that the parameters align with the current image. For the final decision, the cropped view with the highest prediction confidence is selected. Extensive experiments show that IAPL achieves state-of-the-art performance, with mean accuracies of 95.61% and 96.7% on the widely used UniversalFakeDetect and GenImage datasets, respectively. Codes.
@@ 1 Introduction
Driven by the rapid progress in generative AI, particularly GAN [18] and diffusion models [51], generating high-quality synthetic images has become remarkably effortless. These breakthroughs have opened vast opportunities and applications in digital content creation, virtual reality, and personalized media. However, the potential misuse of such technologies raises serious security risks, threatening both public trust and individual privacy [33]. This duality highlights the urgent need for robust and generalizable methods to detect AI-generated images.
@@
Figure 1: Comparison of the learned parameters’ adaptability between our and existing methods when facing a testing image. (a) In the training process, the learned parameters inserted into the CLIP ViT are optimized via a training set that includes limited generative methods. During inference, existing methods (b) keep the learned parameters fixed. In contrast, our method (c) dynamically adjusts the learned parameters based on the test image from unseen sources, enhancing adaptability to the test image.
@@
Many works have been proposed to detect and distinguish AI-generated images from authentic ones. Some methods focus on detecting synthesized faces [34, 62], while others focus on detecting synthesized natural images [68, 61]. Beyond recognizing images generated by seen generators in the training set, a more important task is detecting images produced by unseen generators. The significance lies in improving the model’s generalizability to the ever-evolving generation methods emerging in real-world scenarios. To this end, certain approaches find cues from the low-level traces in the images, such as frequency [53], neighborhood pixel [54], and gradients [52]. However, these methods rely on local details, to some extent overlooking the contribution of unrealistic semantic patterns and overall style in AI-generated content.
@@
Recent state-of-the-art (SOTA) approaches tend to fine-tune visual foundation models, such as CLIP [43], to boost detection performance. These pretrained models encode rich real-world knowledge, which serves as a valuable complement to the limited fake patterns present in training data. Nevertheless, fine-tuned models often struggle to generalize effectively under domain shifts introduced by images from unseen generators, revealing limitations in current adaptation strategies. We contend that achieving stronger generalization requires learnable parameters that remain adaptable to diverse image characteristics. Since images synthesized by different models exhibit substantial variations in texture, semantics, and visual artifacts, relying on fixed model parameters is inadequate for capturing the essential discriminative properties. In this view, we propose a novel Image-Adaptive Prompt Learning (IAPL) method. As shown in Fig. 1, the existing methods keep the parameters learned during training fixed when inferring testing images from unseen sources. In contrast, for IAPL, the learned parameters are dynamically adjusted based on the traits of the image during inference, which enables the model to capture instance-specific cues and better handle variations introduced by different generative methods.
@@
To customize the pre-trained foundation model for the AI-generated image detection task, we first insert some learnable parameters for fine-tuning. Here we combine MLP-based adapters [10] and learnable tokens [66], enabling efficient adaptation while preserving the generalization capability of the backbone model. The former inserts learnable upsampling and downsampling matrices alongside the MLP in the encoder blocks, which are placed at equal intervals throughout the encoder. The latter is applied to the encoder blocks from the 2nd to the $N_{t}$-th. As mentioned above, to further enhance the model’s ability to discriminate images from unseen generators, we introduce the image-adaptive prompt into the first encoder block. It consists of test-time adaptive tokens and conditional information. During training, test-time adaptive tokens are optimized in the same way as standard learnable tokens. At inference, however, they are adjusted based on a single test image, minimizing entropy across different generated views to alleviate the uncertainty introduced by domain shift [50]. The conditional information is handled by the Conditional Information Learner, which employs a convolutional network to extract forgery-specific and image-specific cues from a local rich-texture region. These conditions are instance-dependent, varying across images, and are injected into the detection pipeline to guide feature learning and capture subtle artifacts. Additionally, we introduce a gated mechanism, a control module that adaptively regulates the flow of information. It is employed to connect the test-time adaptive token with the conditional information and propagate the output image-adaptive prompts from the previous to the next layer.
We conduct extensive experiments on two widely used datasets, UniversalFakeDetect [40] and GenImage [68], proving the effectiveness of our method. Specifically, it achieves 95.61% and 96.7% mAcc on these two datasets, respectively, reaching the new SOTA performance. Ablation studies confirm that all proposed components play a role in enhancing performance. As shown in Fig. 2, our method achieves more distinct separation between real and unseen fake images, demonstrating its strong generalization ability. Contributions are as follows:
@@ 2.1 Face Forgery Detection
To ensure the privacy and security of users, face forgery detection has become an important task. Most existing works design sophisticated transformation functions to highlight the fake patterns in generated faces. These transformations function includes frequency-domain transformation [29], spatial-domain operation [59], reconstruction technique [5] and disentanglement [17]. Other works focus on capturing artifacts from facial regions with prominent textures, such as the lips [20] and eyes [32]. For example, LAA-Net [38] targets small, artifact-prone regions and propagates discriminative low-level features into the final output. Nevertheless, as the variety of generative models continues to expand, the effectiveness of methods in handling forgeries from unseen sources remains limited. Thus, some works adopt techniques like blending [49] and adversarial training [8] to enrich the training samples. For example, by modeling and simulating intra-forgery and inter-forgery feature variations in the latent space, LSDA [62] enlarges the forgery space and mitigates the risk of the model overfitting to forgery-specific artifacts.
@@ 2.2 AI-Generated Image Detection
In recent years, generative content has expanded beyond human faces to areas like nature and art. Several approaches capture clues from low-level artifacts, including frequency-based features [23], reconstruction artifacts [60], spatial pixel relationships [54], and gradients [52]. For example, BiHPF [23] leverages two high-pass filters in both the frequency and pixel domains to extract forgery cues. Another important paradigm is to leverage the vision foundation models. These methods benefit from the extensive pre-trained knowledge in visual foundation models and extract high-level forgery features. UniFD [40] is the pioneer to use the CLIP [43] features for forgery detection. FatFormer [35] conducts forgery-aware adapter with frequency analysis to fine-tune CLIP. RINE [28] maps the intermediate features of CLIP to a learnable forgery-aware space through a lightweight network. C2P-CLIP [55] injects category-related concepts to the encoder via contrastive learning. Effort [63] introduces orthogonal subspace decomposition to separate domain-specific and domain-invariant features. Additionally, some works [61, 11] combine low-level and high-level features for more robust representation.
Figure 3: The overview of the image encoder in our method. The trainable parameters added to the original CLIP ViT in our image encoder fall into three categories: MLP-based adapters, learnable tokens, and image-adaptive prompts. MLP-based adapters go into $N_{a}$ encoder blocks at regular intervals. Learnable tokens are used from the $2^{nd}$ to the $N_{t}$-th encoder blocks. Image-adaptive prompts, learned via their dedicated module, are applied as the first block’s input. After training, MLP-based adapters and learnable tokens are fixed, while image-adaptive prompts are adjusted based on the image during inference.
@@ 2.3 Prompt Learning
Prompt learning is initially introduced to substitute handcrafted text templates with learnable tokens [36, 66]. Subsequently, an increasing number of studies began to explore its effectiveness in the visual modality. VPT [24] inserts a small number of learnable tokens into the pretrained ViT [15] layers to adapt the model to downstream tasks. VP [3] uses a learnable image perturbation to be the visual prompt. Then, some multi-modal prompt learning methods are introduced for a visual-language model [27]. Moreover, to improve generalization to broader unseen classes, methods like input-conditional learning and test-time adaptation [57] have been proposed. CoCoOp [65] adds input-conditional tokens to each instance and makes the model less sensitive to class shift. TPT [50] is the pioneer to do Test-Time Tuning in prompt learning by minimizing the averaged entropy to enforce consistent predictions across selected augmentations. R-TPT [48] chooses to minimize pointwise entropy to reduce conflicts.
Several prior studies have explored the application of prompt learning [35, 34] in the context of synthetic image detection. However, the prompts learned by these methods do not generalize well to images generated by unseen generators. In this view, we propose an Image-Adaptive Prompt Learning strategy, which dynamically adjusts the learned prompt based on the characteristics of the test image and increases the adaptability to domain shift.
@@ 3.1 Overview
Figure 4: The overall architecture of Image-Adaptive Prompt Learning. It consists of Test-Time Token Tuning and Conditional Information Learner. The former dynamically adjusts test-time adaptive tokens based on the input image, while the latter extracts conditional information from the texture-rich region, providing forgery-specific and image-specific guidance for prompt generation.
The overview of our image encoder is shown in Fig. 3. We design a flexible yet effective framework that tailors the pre-trained CLIP [43] ViT for AI-generated image detection by integrating three trainable components: MLP-based adapters, learnable tokens, and image-adaptive prompts. These components preserve the backbone’s discriminative power while enabling instance-specific adaptation to improve robustness against unseen generators.
Given an input image, it is first fed into the Image-Adaptive Prompt Learning module to obtain image-adaptive prompts that encode instance-specific cues to guide the backbone model. What’s more, following the vanilla CLIP, the input image is divided into multiple patches and encoded into image tokens $f_{img}\in\mathbb{R}^{N\times D}$, where $N$ and $D$ are the number of tokens and dimension, respectively. After that, the generated image-adaptive prompts, image tokens, and additional CLS token $f_{cls}\in\mathbb{R}^{1\times D}$, are fed into multiple ViT encoder blocks for feature extraction. In each block, the pre-trained parameters of the original CLIP are fixed to preserve knowledge. MLP-based adapters are inserted at equal intervals into $N_{a}$ blocks, while learnable tokens are applied from the $2^{nd}$ to the $N_{t}$-th block. In blocks with learnable tokens, the prompts from the previous block are fused with the tokens via a gated mechanism to form the current input prompts, thereby propagating adaptive information from the image-adaptive prompts and strengthening forgery cue modeling in deeper layers. Finally, the output CLS token from the last encoder block is used for the result prediction via a trainable classifier.
@@ 3.2 Fixed Learned Parameters
Here, we introduce the fixed learned parameters, including adapters and learnable tokens. After training, these parameters remain unchanged and do not adapt to individual input images. They provide a stable backbone that supports the operation of dynamic parameters, ensuring that adaptive adjustments can concentrate on instance-specific forgery cues without undermining the model’s overall generalization.
MLP-based Adapter. Unlike other methods that use LoRA [22, 55] or carefully designed adapters [35], we employ a simple yet effective MLP-based adapter [10]. This design aims to minimize the computational cost associated with adapters, allowing more resources to be dedicated to the Image-Adaptive Prompt Learning module introduced subsequently. As shown in Fig. 3, the MLP-based adapter includes a down-projection $W_{d}\in\mathbb{R}^{D\times\hat{D}}$ and a up-projection $W_{d}\in\mathbb{R}^{\hat{D}\times D}$, where $\hat{D}$ is the middle dimension and satisfies $\hat{D}\ll D$. A ReLU activation function and dropout technique are employed between the up-projection and down-projection to enhance non-linearity and prevent overfitting. The adapter’s output is scaled by a factor $s$ and subsequently added to the original output. For parameter initialization, we follow the strategy of LoRA [22], the down-projection is initialized with Kaiming uniform [21], while the up-projection is initialized to zeros.
Learnable Token. Formally, in the $j$-th block $(1<j\leq N_{t})$, a set of learnable tokens $T^{(j)}\in\mathbb{R}^{M\times D}$ is fused with the output prompts $P_{o}^{(j-1)}\in\mathbb{R}^{M\times D}$ from the former block through a gated mechanism to produce the input prompts $P_{i}^{(j)}\in\mathbb{R}^{M\times D}$ of this block. Here, $M$ denotes the number of learnable tokens, and $D$ is the embedding dimension. The gated mechanism is defined in Eq. 1.
@@
where $\alpha^{(j)}\in\mathbb{R}^{D}$ denotes the channel-wise gating coefficients in the $j$-th layer. Initialized to a small value (e.g., $1\times 10^{-6}$) and optimized during training, these coefficients are applied directly to the prompt, thereby controlling on a per-channel basis how the fused prompt incorporates information from the former layer. After that, the generated prompts $P_{i}^{(j)}$ are concatenated with the image and CLS tokens, and then fed into the subsequent encoder blocks for further representation learning.

### v3

@@ Abstract
In AI-generated image detection, current cutting-edge methods typically adapt pre-trained foundation models through partial-parameter fine-tuning. However, these approaches often struggle to generalize to forgeries from unseen generators, as the fine-tuned models capture only limited patterns from training data and fail to reflect the evolving traits of new ones. To overcome this limitation, we propose Image-Adaptive Prompt Learning (IAPL), a novel paradigm that dynamically adjusts the prompts fed into the encoder according to each testing image, rather than fixing them after training. This design significantly enhances robustness and adaptability to diverse forged images. The dynamic prompts integrate conditional information with test-time adaptive tokens through a lightweight learnable scaling factor. The conditional information is produced by a Conditional Information Learner, which leverages CNN-based feature extractors to model both forgery-specific and general conditions. The test-time adaptive tokens are optimized during inference on a single sample by enforcing prediction consistency across multiple views, ensuring that the parameters align with the current image. For the final decision, the optimal input with the highest prediction confidence is selected. Extensive experiments show that IAPL achieves state-of-the-art performance, with mean accuracies of 95.61% and 96.7% on the widely used UniversalFakeDetect and GenImage datasets, respectively. Codes.
@@ 1 Introduction
Driven by the rapid progress in generative AI, particularly GAN [goodfellow2020generative] and diffusion models [song2020score], generating high-quality synthetic images has become remarkably effortless. These breakthroughs have opened vast opportunities and applications in digital content creation, virtual reality, and personalized media. However, the potential misuse of such technologies raises serious security risks, threatening both public trust and individual privacy [li2025unleashing]. This duality highlights the urgent need for robust and generalizable methods to detect AI-generated images.
@@
Figure 1: Comparison of the adaptability of prompts between our and existing methods when facing a testing image. (a) In the training process, the parameters of learnable prompts are optimized via a training set that includes limited generative methods. During inference, existing methods (b) keep the learned prompts fixed. In contrast, our method (c) dynamically adjusts the learned prompts based on test images from unseen sources, thereby enhancing generalization to novel test images.
@@
Several works have been proposed to detect and distinguish between AI-generated images and authentic ones. Some methods focus on detecting synthesized faces [lin2025standing, yan2024transcending], while others focus on detecting synthesized natural images [zhu2023genImage, yan2024sanity]. Beyond recognizing images generated by seen generators in the training set, a more important task is detecting images produced by unseen generators. The significance lies in improving the model’s generalizability to the ever-evolving generation methods emerging in real-world scenarios. To this end, certain approaches find cues from the low-level traces in the images, such as frequency [tan2024frequency], neighborhood pixel [tan2024rethinking], and gradients [tan2023learning]. However, these methods rely on local details, to some extent overlooking the contribution of unrealistic semantic patterns and overall style in AI-generated content.
@@
Recent state-of-the-art (SOTA) approaches tend to fine-tune visual foundation models, such as CLIP [radford2021learning], to boost detection performance. These pre-trained models encode rich real-world knowledge, which serves as a valuable complement to the limited fake patterns present in training data. Nevertheless, fine-tuned models often struggle to generalize effectively under domain shifts introduced by images from unseen generators, revealing limitations in current adaptation strategies. We contend that achieving stronger generalization requires learnable parameters that remain adaptable to diverse image characteristics. Since images synthesized by different models exhibit substantial variations in texture, semantics, and visual artifacts, relying on fixed model parameters is inadequate for capturing the essential discriminative properties. In this view, we propose a novel Image-Adaptive Prompt Learning (IAPL) method. As shown in Fig. 1, the existing methods [shi2025mirage, lin2025standing] keep the prompts learned during training fixed when inferring testing images from unseen sources. In contrast, for IAPL, the prompts input to the image encoder are dynamically adjusted based on the traits of the image during inference, which enables the model to capture instance-specific cues and better handle variations introduced by different generative methods.
@@
The image-adaptive prompts are input into the first encoder block and consist of test-time adaptive tokens and conditional information. During training, test-time adaptive tokens are optimized in the same way as standard learnable tokens. At inference, however, they are adjusted based on a single test image, minimizing entropy across different generated views to alleviate the uncertainty introduced by domain shift [shu2022test]. The conditional information is processed by the Conditional Information Learner, which leverages high-pass filters and CNN-based feature extractors to capture both forgery-specific and general cues from local regions rich in texture. These conditions are instance-dependent, varying across images, and are injected into the detection pipeline to guide feature learning and capture subtle artifacts. To better fuse test-time adaptive token and conditional information, we further introduce the learnable scaling factor, a control module that adaptively regulates information flow. It aggregates the two types of information using learnable coefficients, emphasizing relevant signals and suppressing less informative ones. Moreover, to equip the pre-trained foundation model with enhanced task-specific capabilities for AI-generated image detection, we also insert some learnable parameters for fine-tuning, which will be fixed after training. Here, we combine MLP-based adapters [chen2022adaptformer] and learnable tokens [zhou2022learning], enabling efficient adaptation while preserving the generalization capability of the backbone model. The former inserts learnable upsampling and downsampling matrices alongside the MLP module in the encoder blocks, which are placed at equal intervals throughout the encoder. The latter is applied to the encoder blocks from the 2nd to the $N_{t}$-th.
We conduct extensive experiments on two widely used datasets, UniversalFakeDetect [ojha2023towards] and GenImage [zhu2023genImage], proving the effectiveness of our method. Specifically, it achieves 95.61% and 96.7% mAcc on these two datasets, respectively, reaching the new SOTA performance. Ablation studies confirm that all proposed components play a role in enhancing performance. As shown in Fig. 2, our method achieves more distinct separation between real and unseen fake images, demonstrating its strong generalization ability. Contributions are as follows:
@@ 2.1 Face Forgery Detection
To ensure the privacy and security of users, face forgery detection has become an important task. Most existing works design sophisticated transformation functions to highlight the fake patterns in generated faces. These transformations function includes frequency-domain transformation [li2021frequency, frank2020leveraging], spatial-domain operation [wang2023noise], reconstruction technique [cao2022end] and disentanglement [fu2025exploring]. Other works focus on capturing artifacts from facial regions with prominent textures, such as the lips [haliassos2021lips] and eyes [li2018ictu]. For example, LAA-Net [nguyen2024laa] targets small, artifact-prone regions and propagates discriminative low-level features into the final output. Nevertheless, as the variety of generative models continues to expand, the effectiveness of methods in handling forgeries from unseen sources remains limited. Thus, some works adopt techniques like blending [shiohara2022detecting] and adversarial training [chen2022self] to enrich the training samples. For example, by modeling and simulating intra-forgery and inter-forgery feature variations in the latent space, LSDA [yan2024transcending] enlarges the forgery space and mitigates the risk of the model overfitting to forgery-specific artifacts.
@@ 2.2 AI-Generated Image Detection
In recent years, generative content has expanded beyond human faces to areas like nature and art. Several approaches capture clues from low-level artifacts, including frequency-based features [jeong2022bihpf], reconstruction artifacts [wang2023dire], spatial pixel relationships [tan2024rethinking], and gradients [tan2023learning]. For example, BiHPF [jeong2022bihpf] leverages two high-pass filters in both the frequency and pixel domains to extract forgery cues. Another important paradigm is to leverage the vision foundation models. These methods benefit from the extensive pre-trained knowledge in visual foundation models and extract high-level forgery features. UniFD [ojha2023towards] is the pioneer to use the CLIP [radford2021learning] features for forgery detection. FatFormer [liu2024forgery] conducts forgery-aware adapter with frequency analysis to fine-tune CLIP. RINE [koutlis2024leveraging] maps the intermediate features of CLIP to a learnable forgery-aware space through a lightweight network. C2P-CLIP [tan2025c2p] injects category-related concepts to the encoder via contrastive learning. Effort [yan2024orthogonal] introduces orthogonal subspace decomposition to separate domain-specific and domain-invariant features. Additionally, some works [yan2024sanity, cheng2025co] combine low-level and high-level features for more robust representation.
Figure 3: Overview of the detection pipeline in our method. The trainable parameters added to the original CLIP ViT in our image encoder fall into three categories: MLP-based adapters, learnable tokens, and image-adaptive prompts. MLP-based adapters go into $N_{a}$ encoder blocks at regular intervals. Learnable tokens are used from the $2^{nd}$ to the $N_{t}$-th encoder blocks. Image-adaptive prompts, learned via their dedicated module, are applied as the first block’s input. After training, MLP-based adapters and learnable tokens are fixed, while image-adaptive prompts are adjusted based on the image during inference.
@@ 2.3 Prompt Learning
Prompt learning is initially introduced to substitute handcrafted text templates with learnable tokens [liu2021p, zhou2022learning]. Subsequently, an increasing number of studies began to explore its effectiveness in the visual modality. VPT [jia2022visual] inserts a small number of learnable tokens into the pretrained ViT [dosovitskiy2020image] layers to adapt the model to downstream tasks. VP [bahng2022exploring] uses a learnable image perturbation to be the visual prompt. Then, some multi-modal prompt learning methods are introduced for a visual-language model [khattak2023maple]. Moreover, to improve generalization to broader unseen classes, methods like input-conditional learning and test-time adaptation [wang2020tent] have been proposed. CoCoOp [zhou2022conditional] adds input-conditional tokens to each instance and makes the model less sensitive to class shift. TPT [shu2022test] is the pioneer to do Test-Time Tuning in prompt learning by minimizing the averaged entropy to enforce consistent predictions across selected augmentations. R-TPT [sheng2025r] chooses to minimize pointwise entropy to reduce conflicts.
Several prior studies have explored the application of prompt learning [liu2024forgery, lin2025standing] in the context of synthetic image detection. However, the prompts learned by these methods do not generalize well to images generated by unseen generators. In this view, we propose an Image-Adaptive Prompt Learning strategy, which dynamically adjusts the learned prompt based on the characteristics of the test image and increases the adaptability to domain shift.
@@ 3.1 Overview
Figure 4: The overall architecture of Image-Adaptive Prompt Learning. It consists of Test-Time Token Tuning and Conditional Information Learner. The former dynamically adjusts test-time adaptive tokens based on the input image, while the latter extracts conditional information from the texture-rich region, providing forgery-specific and general guidance for prompt generation.
The overview of our detection pipeline is shown in Fig. 3. We design a flexible yet effective framework that tailors the pre-trained CLIP [radford2021learning] ViT for AI-generated image detection by integrating three trainable components: MLP-based adapters, learnable tokens, and image-adaptive prompts. These components preserve the backbone’s discriminative power while enabling instance-specific adaptation to improve robustness against unseen generators.
Given an input image, it is first fed into the Image-Adaptive Prompt Learning module to obtain image-adaptive prompts that encode image-oriented cues to guide the backbone model. What’s more, following the vanilla CLIP, the input image is divided into multiple patches and encoded into image tokens $f_{img}\in\mathbb{R}^{N\times D}$, where $N$ and $D$ are the number of tokens and dimension, respectively. After that, the generated image-adaptive prompts, image tokens, and additional CLS token $f_{cls}\in\mathbb{R}^{1\times D}$, are fed into multiple encoder blocks for feature extraction. In each block, the pre-trained parameters of the original CLIP are fixed to preserve knowledge. MLP-based adapters are inserted at equal intervals into $N_{a}$ blocks, while learnable tokens are applied from the $2^{nd}$ to the $N_{t}$-th block. In blocks with learnable tokens, the prompts from the previous block are fused with the tokens via a learnable scaling factor to form the current input prompts, thereby propagating adaptive information from the image-adaptive prompts and strengthening forgery cue modeling in deeper layers. Finally, the output CLS token from the last encoder block is used for the result prediction via a trainable classifier.
@@ 3.2 Fixed Learned Parameters
We first introduce the MLP-based adapters and learnable tokens, which remain unchanged after training and do not vary with input images. Therefore, they are referred to as fixed learned parameters. It is worth noting that, although our paper emphasizes the importance of dynamic parameters in capturing instance-specific forgery cues, the fixed learned parameters still play a crucial role: they provide a stable backbone for the model, allowing dynamic parameters to focus on adaptive adjustments for each input image while preventing excessive flexibility that could compromise overall inference accuracy.
MLP-based Adapter. As shown in Fig. 3, we employ a simple yet effective MLP-based adapter [chen2022adaptformer]. It includes a down-projection $W_{d}\in\mathbb{R}^{D\times\hat{D}}$ and a up-projection $W_{d}\in\mathbb{R}^{\hat{D}\times D}$, where $\hat{D}$ is the middle dimension and satisfies $\hat{D}\ll D$. A ReLU activation function and dropout technique are employed between the up-projection and down-projection to enhance non-linearity and prevent overfitting. The adapter’s output is scaled by a factor $s$ and subsequently added to the original output. For parameter initialization, we follow LoRA [hu2022lora]: the down-projection is initialized using Kaiming uniform initialization [he2015delving], while the up-projection is initialized to zeros.
Learnable Token. Formally, in the $j$-th block $(2\leq j\leq N_{t})$, a set of learnable tokens $T^{(j)}\in\mathbb{R}^{M\times D}$ is fused with the output prompts $P_{o}^{(j-1)}\in\mathbb{R}^{M\times D}$ from the former block through a learnable scaling factor to produce the input prompts $P_{i}^{(j)}\in\mathbb{R}^{M\times D}$ of this block. Here, $M$ denotes the number of learnable tokens, and $D$ is the embedding dimension. The learnable scaling factor is defined in Eq. 1.
@@
where $\alpha^{(j)}\in\mathbb{R}^{D}$ denotes the channel-wise coefficients in the $j$-th layer. These coefficients are initialized to small values and optimized during training. They are applied directly to the prompt, enabling fine-grained control over how information from the previous layer is integrated on a per-channel basis. In other words, $\alpha^{(j)}$ modulates the contribution of each channel in the fused prompt, allowing the model to emphasize or suppress specific features dynamically. Then, the generated prompts $P_{i}^{(j)}$ are concatenated with the image and CLS tokens and fed into the subsequent encoder blocks for further representation learning.